OpenStudy (anonymous):

4/9 - 1/3.. Im learning this..can u spleases show me the steps

OpenStudy (anonymous):

You need a common denominator: 4/9 - 1/3 4/9 - 3/9 = 1/9

OpenStudy (anonymous):

a/b+c/d = (ad+bc)/bd always